OpenStudy (anonymous):

The number of seven lettered words that can be formed using all the letters of the word FACULTY so that the words end in F and begins with Y is?

OpenStudy (anonymous):

ignore the F and Y because you have no choice there, so answer is \(5!\)

OpenStudy (anonymous):

the answer is 120

OpenStudy (anonymous):

yes, \(5!=5\times 4\times 3\times 2=120\)

OpenStudy (anonymous):

@ajprincess @dumbcow @Mertsj

OpenStudy (anonymous):

@TuringTest @blues

OpenStudy (anonymous):

@hoblos@apoorvk plzz help

OpenStudy (anonymous):

help plz

OpenStudy (apoorvk):

 so F and Y are already fixed, you just have to find all 5 digit- combinations using the left over 5 letters, that will fill in between. helps?